If X has a Standard Gamma distribution with α=7.

  • Compute E(X)=?
  • Compute σx=?
  • Compute P(X≤5)=?
  • Compute P(X>8)=?
  • Compute P(3<X<8)=?
Answer #1

as for standard gamma distribution: =1

a)

E(X)= α =7*1 =7

b) σx= sqrt(α 2)=sqrt(7*12)=2.6458

c)

P(X<=5)=0.237817

d)

P(X>8)=0.313374

e)

P(3<X<8)=0.653117